. Student dissolves a tablet of Vitamin C in 100.0 mL of water. 2. Using 5.00 mL volumetric pipet, student transfers 5.00 mL of the tablet solution in

to an Erlenmeyer flask. 3. After adding starch indicator, student titrated Vitamin C containing solution with standard solution of Iodine (5.00E-3 M). It took 17.73 mL of Iodine solution. MM of vitamin c = 176.1 g/mol Calculate mass of Vitamin C in the tablet.
Answer #1

Number of mole of Iodine solution = Volume of solution (in L) * Molarity (M) = 17.73/1000 * 5 * 10^(-3) = 8.865 * 10^(-5) moles

So, number of moles of Vitamin C present in 5.00 ml pipet will be 8.865 * 10^(-5) moles

Number of Vitamin C dissolved in 100.0 ml of water = 8.865 * 10^(-5) * 20 = 1.773 * 10^(-3) moles

Mass of Vitamin C in tablet = (Number of moles of Vitamin C) * (Molar mass of Vitamin C) = (1.773 * 10^(-3) mol) * (176.1 g/mol) = 0.312 grams

Note - Post any doubts/queries in comments section.